# Unified Thinker: A General Reasoning Modular Core for Image Generation

### v1

@@ Appendix C Details of HieraReason-40K
HieraReason-40K is built to train a generator-agnostic Thinker that produces structured reasoning traces and a final enhanced prompt for downstream diffusion models. We collect 40K instruction examples from four open-source datasets covering image editing and reasoning-oriented generation/editing tasks han2025unireditbench; huang2025interleaving; qian2025pico; fang2025fluxreason. Specifically, we sample 10K instances from each source, and then convert them into a unified format via structured knowledge distillation with Gemini under our system prompt, yielding intermediate reasoning traces aligned with the final enhanced prompt.

### v2

@@ Appendix C Details of HieraReason-40K
HieraReason-40K is built to train a generator-agnostic Thinker that produces structured reasoning traces and a final enhanced prompt for downstream diffusion models. We collect 40K instruction examples from four open-source datasets covering image editing and reasoning-oriented generation/editing tasks Han et al. (2025); Huang et al. (2025); Qian et al. (2025); Fang et al. (2025). Specifically, we sample 10K instances from each source, and then convert them into a unified format via structured knowledge distillation with Gemini under our system prompt, yielding intermediate reasoning traces aligned with the final enhanced prompt.